OpenStudy (anonymous):

find the real solutions if any x^2+5x+1=0

jimthompson5910 (jim_thompson5910):

use the quadratic formula to solve for x \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(5)\pm\sqrt{(5)^2-4(1)(1)}}{2(1)}\] I'll let you finish up